# Downlink and Uplink ISAC in Continuous-Aperture Array (CAPA) Systems

### v1

@@ Abstract
A continuous-aperture array (CAPA)-based integrated sensing and communications (ISAC) framework is proposed for both downlink and uplink scenarios. Within this framework, continuous operator-based signal models are employed to describe the sensing and communication processes. The performance of communication and sensing is analyzed using two information-theoretic metrics: the communication rate (CR) and the sensing rate (SR). 1) For downlink ISAC, three continuous beamforming designs are proposed: i) the communications-centric (C-C) design that maximizes the CR, ii) the sensing-centric (S-C) design that maximizes the SR, and iii) the Pareto-optimal design that characterizes the Pareto boundary of the CR-SR region. A signal subspace-based approach is proposed to derive the closed-form optimal beamformers for the considered designs. On this basis, closed-form expressions are derived for the achievable CRs and SRs, and the downlink rate region achieved by CAPAs is characterized. 2) For uplink ISAC, the C-C and S-C successive interference cancellation (SIC)-based methods are proposed to manage inter-functionality interference. Using the subspace approach along with the time-sharing technique, closed-form expressions for the optimal beamformers are derived, and the achievable CRs, SRs, and rate region are analyzed. Numerical results demonstrate that, for both downlink and uplink, CAPA-based ISAC achieves higher CRs and SRs as well as larger CR-SR regions compared to conventional spatially discrete array (SPDA)-based ISAC.
@@ I Introduction
Among the key research topics in ISAC, a significant focus is the integration of multiple-antenna technologies. Specifically, technologies such as multiple-input multiple-output (MIMO) provide additional spatial degrees of freedom (DoFs) and enhanced array gains for both communications and sensing. These improvements lead to higher network throughput and better sensing resolution; see [2, 4] and related references for more details. In general, the available spatial DoFs and array gains scale linearly with the number of antennas [5]. Motivated by this, various evolved large-scale MIMO variants and novel array architectures have been proposed and applied to ISAC. Examples include massive MIMO [6], holographic MIMO [7], gigantic MIMO [8], extremely large aperture arrays (ELAAs) [9], reconfigurable intelligent surfaces (RISs) [10], and fluid/movable antennas [11, 12]. These works provide extensive results demonstrating the significant performance gains in ISAC systems achieved by leveraging these novel multiple-antenna technologies.
@@
Despite their diverse architectures, these novel multiple-antenna systems share a common evolutionary trend: larger aperture sizes, denser antenna configurations, higher operating frequencies, and more flexible structures. This trend leads to the development of an (approximately) continuous electromagnetic (EM) aperture, also referred to as a continuous-aperture array (CAPA) [13]. A CAPA operates as a single large-aperture antenna with a continuous current distribution, comprising a (virtually) infinite number of radiating elements coupled with electronic circuits and fed by a limited number of radio-frequency (RF) chains [13]. Compared with conventional spatially discrete arrays (SPDAs), a CAPA fully utilizes the entire surface of the aperture and enables free control of the current distribution. This capability yields significantly enhanced spatial DoFs and array gains. Initial efforts to apply CAPAs to both communications and sensing have demonstrated that CAPA outperforms traditional SPDAs in both areas, as discussed in Section I-A.
@@
Given this context, it is natural to explore CAPA technology to further enhance ISAC performance. However, due to its continuous nature, a CAPA-based system should be modeled using EM theory and based on a continuous Hilbert–Schmidt operator-based integral model, which differs significantly from the conventional matrix-based model for SPDAs [14, 13, 15, 16, 17, 18]. This shift renders existing SPDA-based ISAC frameworks inapplicable to CAPA-based ISAC systems. Therefore, a new conceptual CAPA-based ISAC framework is needed to address the continuous EM nature inherent in such systems.
I-A Prior Works
@@
Recently, CAPA has attracted growing research interest. The performance of CAPA-based communications has been extensively studied in terms of spatial DoFs [19, 20, 21], array gains [22, 23], and channel capacity [24], with most of these studies focusing on line-of-sight (LoS) channels. To address this limitation, researchers have developed a novel multipath spatial fading model for CAPA [18]. Using this model, the work in [25] analyzed the spatial correlation and the diversity-multiplexing trade-off in CAPA-based communications. These studies have collectively validated CAPA’s effectiveness in enhancing spatial DoFs and array gains, thereby providing a theoretical foundation for its application.
@@
Another key research area is continuous beamforming design for CAPA-based communications. Since optimizing a continuous source current function involves challenging fractional programming, various approaches have been proposed. For example, a discretization-based approach (using Fourier series expansion) has been proposed to approximate continuous signals, which simplified the optimization problem [26, 27]. Additionally, a calculus of variations method was introduced to design low-complexity beamforming [28] and explore optimal beamforming structures [29, 30].
@@ I-B Motivation and Contributions
Motivated by the research gaps discussed above, this article presents a comprehensive performance analysis of a CAPA-based ISAC system. We analyze the SRs, CRs, and rate regions for both downlink and uplink scenarios. The main contributions are summarized as follows.
@@
  • For downlink ISAC, we propose three continuous DFSAC beamforming designs: i) a communications-centric (C-C) design that maximizes the CR, ii) a sensing-centric (S-C) design that maximizes the SR, and iii) a Pareto-optimal design that characterizes the Pareto boundary of the rate region. For both C-C and S-C designs, we derive closed-form expressions for SRs and CRs. For the Pareto-optimal design, we propose a signal subspace-based approach to derive closed-form solutions for the optimal beamformer and the achievable rate region.
@@
  • We present numerical results to demonstrate that: i) CAPA-based ISAC achieves higher SRs and CRs than both SPDA-based ISAC and CAPA-based FDSAC in both downlink and uplink scenarios; ii) The achievable SR-CR rate regions of SPDA-based ISAC and CAPA-based FDSAC systems are entirely encompassed within those of the CAPA-based ISAC system.
@@ II-B1 Communication Model
For LoS scenarios, the channel response $h_{\rm{c}}(\mathbf{u},{\mathbf{t}})$ is expressed as follows [23, 22]:
@@
where $\sqrt{\frac{\lvert{\mathbf{e}}_{\rm{t}}^{\mathsf{T}}(\mathbf{u}-\mathbf{t})\rvert}{\lVert\mathbf{u}-\mathbf{t}\rVert}}$ accounts for the effective aperture loss caused by misalignment between the signal propagation direction and the array’s normal direction. Moreover, ${\mathbf{e}}_{y}=[0,1,0]^{\mathsf{T}}$ denotes the normal vector to the transmit CAPA, and $g(\cdot,\cdot)$ is the scalar Green’s function that models the effects of free-space electromagnetic propagation. By [23, 22], we have
$g(\mathbf{u},\mathbf{t})=-{\rm{j}}\eta k_{0}\frac{{\rm{e}}^{-{\rm{j}}k_{0}\lVert\mathbf{u}-\mathbf{t}\rVert}}{4\pi\lVert\mathbf{u}-\mathbf{t}\rVert}\left(1+\frac{{\rm{j}}k_{0}^{-1}}{\lVert\mathbf{u}-\mathbf{t}\rVert}-\frac{k_{0}^{-2}}{\lVert\mathbf{u}-\mathbf{t}\rVert^{2}}\right),$ (9)
@@
$g(\mathbf{u},\mathbf{t})=-{\rm{j}}\eta k_{0}\frac{{\rm{e}}^{-{\rm{j}}k_{0}\lVert\mathbf{u}-\mathbf{t}\rVert}}{4\pi\lVert\mathbf{u}-\mathbf{t}\rVert}.$ (10)
@@ II-B2 Sensing Model
where $h_{\rm{s}}(\mathbf{r},{\mathbf{t}})$ represents the target response, and $n_{{\rm{s}},\ell}(\mathbf{r})$ is additive noise satisfying
@@
In this work, we assume that the target’s location is perfectly tracked and focus on estimating the RCS $\beta_{\rm{s}}$. Under this scenario, the sensing task aims to extract the environmental information contained in $\beta_{\rm{s}}$ from the received echo signals $\{y_{{\rm{s}},\ell}(\mathbf{r})\}_{\ell=1}^{L}$, given knowledge of ${\mathbf{x}}({\mathbf{t}})$. According to the Swerling-I model [38], the RCS $\beta_{\rm{s}}$ is assumed to remain relatively constant across pulses and has an a prior Rayleigh-distributed amplitude. Under this model, $\beta_{\rm{s}}$ follows a complex Gaussian distribution, i.e., $\beta_{\rm{s}}\sim{\mathcal{CN}}(0,\alpha_{\rm{s}})$, where $\alpha_{\rm{s}}>0$ is the arithmetic mean of all RCS values of the reflecting object. This parameter represents the average reflection strength [38].
@@ II-C Uplink ISAC
$x_{{\rm{c}},\ell}({\mathbf{u}})=\sqrt{P_{\rm{c}}}w_{\rm{u}}({\mathbf{u}})s_{{\rm{c}},\ell},$ (20)
@@ III-A2 Performance Metric for Sensing
The next task is to recover $\beta_{\rm{s}}$ from $\hat{\mathbf{y}}_{{\rm{s}}}$ given ${\mathbf{s}}$. The information-theoretic limits of this sensing task are measured by the sensing mutual information (MI), which quantifies the MI between $\hat{\mathbf{y}}_{{\rm{s}}}$ and $\beta_{\rm{s}}$, conditioned on the ISAC signal ${\mathbf{s}}$ [35]. To evaluate sensing performance, we use the sensing rate (SR) as the performance metric, which is defined as the sensing MI per unit time. Since each DFSAC symbol lasts one unit of time, the SR is expressed as follows:
@@ III-B2 Performance of Communications
Please refer to Appendix -A for more details. ∎
@@ III-B3 Performance of Sensing
$\begin{split}\rho_{\mathrm{d}}=&\frac{\pi^{2}L_{x}L_{z}}{4N^{2}}\sum\nolimits_{n=1}^{N}{\sum\nolimits_{n^{\prime}=1}^{N}{\sqrt{\left(1-\xi_{n}^{2}\right)(1-\xi_{n^{\prime}}^{2})}}}\\ &\times\hat{h}_{\mathrm{c}}^{*}\Big{(}\frac{\xi_{n}+1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)}\hat{h}_{\mathrm{s}}\Big{(}\frac{\xi_{n}+1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)},\end{split}$ (44)
@@
$\hat{h}_{k}(x,z)\triangleq\frac{\mathrm{j}\eta k_{0}\sqrt{\frac{r_{k}\Psi_{k}}{4\pi}}\mathrm{e}^{-\mathrm{j}k_{0}(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{\frac{1}{2}}}}{(x^{2}\!+\!z^{2}\!-\!2r_{k}\!\left(\Phi_{k}x\!+\!\Theta_{k}z\right)\!+\!r_{k}^{2})^{\frac{3}{4}}}$ (45)
@@
The term $\int_{\mathcal{A}_{\mathrm{r}}}{\left|a_{\mathrm{r}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}=g_{\mathrm{r}}$ can be computed by following steps similar to those in Appendix -A. Additionally, the term $\rho_{\mathrm{d}}$ can be expressed as follows:
$\rho_{\mathrm{d}}=\int_{-{L_{z}}/{2}}^{{L_{z}}/{2}}{\int_{0}^{L_{x}}{\hat{h}_{\mathrm{c}}^{*}(x,z)\hat{h}_{\mathrm{s}}(x,z)\mathrm{d}x\mathrm{d}z}}.$ (46)

### v2

@@ Abstract
A continuous-aperture array (CAPA)-based integrated sensing and communications (ISAC) framework is proposed for both downlink and uplink scenarios. Within this framework, continuous operator-based signal models are employed to describe the sensing and communication processes. The performance of communication and sensing is analyzed using two information-theoretic metrics: the communication rate (CR) and the sensing rate (SR). 1) For downlink ISAC, three continuous beamforming designs are proposed: i) the communications-centric (C-C) design that maximizes the CR, ii) the sensing-centric (S-C) design that maximizes the SR, and iii) the Pareto-optimal design that characterizes the Pareto boundary of the CR-SR region. A low-complexity signal subspace-based approach is proposed to derive the closed-form optimal beamformers for the considered designs. On this basis, closed-form expressions are derived for the achievable CRs and SRs, and the downlink rate region achieved by CAPAs is characterized. 2) For uplink ISAC, the C-C and S-C successive interference cancellation-based methods are proposed to manage inter-functionality interference. Using the subspace approach closed-form expressions for the optimal detectors as well as the achievable CRs and SRs are derived. The uplink SR-CR region is characterized based on the time-sharing technique. Numerical results demonstrate that, for both downlink and uplink, CAPA-based ISAC achieves higher CRs and SRs as well as larger CR-SR regions compared to conventional spatially discrete array-based ISAC.
@@ I Introduction
Among the key research topics in ISAC, a significant focus is the integration of multiple-antenna technologies. Specifically, technologies such as multiple-input multiple-output (MIMO) provide additional spatial degrees of freedom (DoFs) and enhanced array gains for both communications and sensing. These improvements lead to higher network throughput and better sensing resolution; see [2, 4] and related references for more details. In general, the available spatial DoFs and array gains scale linearly with the number of antennas. Motivated by this, various evolved large-scale MIMO variants and novel array architectures have been proposed and applied to ISAC. Examples include massive MIMO [5], holographic MIMO [6], extremely large aperture arrays (ELAAs) [7], reconfigurable intelligent surfaces (RISs) [8], and fluid/movable antennas [9, 10]. These works provide extensive results demonstrating the significant performance gains in ISAC systems achieved by leveraging these novel multiple-antenna technologies.
@@
Despite their diverse architectures, these novel multiple-antenna systems share a common evolutionary trend: larger aperture sizes, denser antenna configurations, higher operating frequencies, and more flexible structures. This trend leads to the development of an (approximately) continuous electromagnetic (EM) aperture, also referred to as a continuous-aperture array (CAPA) [11]. A CAPA operates as a single large-aperture antenna with a continuous current distribution, comprising a (virtually) infinite number of radiating elements coupled with electronic circuits and fed by a limited number of radio-frequency (RF) chains [11]. Compared with conventional spatially discrete arrays (SPDAs), a CAPA fully utilizes the entire surface of the aperture and enables free control of the current distribution. This capability yields significantly enhanced spatial DoFs and array gains. Initial efforts to apply CAPAs to both communications and sensing have demonstrated that CAPA outperforms traditional SPDAs in both areas, as discussed in Section I-B.
@@
Given this context, it is natural to explore CAPA technology to further enhance ISAC performance. However, due to its continuous nature, a CAPA-based system should be modeled using EM theory and based on a continuous Hilbert–Schmidt operator-based integral model [12, 11, 13, 14, 15, 16], which differs significantly from the conventional matrix-based model for SPDAs. This shift renders existing SPDA-based ISAC frameworks inapplicable to CAPA-based ISAC systems. Therefore, a new conceptual CAPA-based ISAC framework is needed to address the continuous EM nature inherent in such systems.
I-A Hardware Implementation of CAPAs
Achieving a continuous-aperture array (CAPA) has been a longstanding goal in antenna design, with early contributions from Harold Wheeler [17] and David Staiman [18] in the 1960s, who laid the groundwork for this design ideal. Recently, advancements in EM and optical materials and array fabrication have enabled the development of CAPA prototypes, utilizing designs such as metasurface-based leaky-wave antennas [19, 20], optically driven tightly coupled arrays [21], and interdigital transducer (IDT)-based grating antennas [22]. Some prototypes have even reached commercialization and, according to initial test reports, demonstrate significant potential for enhancing coverage and throughput in practical wireless network environments [23, 24].
Although these implementations differ, they share a common objective: to realize CAPA as an architecture for analog EM beamforming, in contrast to traditional digital beamforming approaches [11]. Fundamentally, CAPA is characterized not only by a continuous or semi-continuous EM structure but more critically by its ability to support full analog control of both amplitude and phase across the aperture [25, 11]. This enables CAPA systems to require only as many RF chains as there are spatially multiplexed data streams. In contrast, conventional SPDAs lack reconfigurability at the antenna level—their radiation patterns are fixed once the antenna type is selected. As a result, beamforming in SPDAs must rely entirely on analog/digital processing before the RF front end.
However, achieving full analog amplitude and phase control over a continuous aperture at the radio front end introduces significant hardware complexity. This remains one of the principal challenges in CAPA implementation. A detailed discussion of the associated design considerations and enabling technologies is provided in [11].
Additionally, despite its promise for full analog EM beamforming, CAPA faces hardware impairments similar to those of conventional arrays—including component non-linearity, I/Q imbalance, quantization error, and phase noise. Additional concerns, such as EM coupling, feed network complexity, and fabrication constraints, must also be addressed [26]. These challenges suggest that even with current advancements, the realization of a fully functional CAPA or densely packed semi-CAPA remains a significant undertaking, requiring collaborative innovation from both the communication and antenna communities.
I-B Prior Works
@@
While the hardware implementation of an ideal CAPA is still at an early stage, growing attention has been directed toward its theoretical analysis and designs of CAPA-based communications and sensing. The performance of CAPA-based communications has been extensively studied in terms of spatial DoFs [27], array gains [28], and channel capacity [26], with most of these studies focusing on line-of-sight (LoS) channels. To address this limitation, researchers have developed a novel multipath spatial fading model for CAPA [16]. Using this model, the work in [29] analyzed the spatial correlation and the diversity-multiplexing trade-off in CAPA-based communications. These studies have collectively validated CAPA’s effectiveness in enhancing spatial DoFs and array gains, thereby providing a theoretical foundation for its application.
@@
Another key research area is continuous beamforming design for CAPA-based communications. Since optimizing a continuous source current function involves challenging fractional programming, various approaches have been proposed. For example, a discretization-based approach (using Fourier series expansion) has been proposed to approximate continuous signals, which simplified the optimization problem [30, 31, 32]. Additionally, a calculus of variations method was introduced to design low-complexity beamforming [33] and explore optimal multi-user beamforming structures [34].
@@ I-C Motivation and Contributions
Motivated by the above research gaps, this article presents a comprehensive performance analysis of a CAPA-based ISAC system. We propose a signal subspace-based approach with significantly lower complexity than the widely used Fourier-based approach [30, 31, 32, 38], and employ it to study the SRs, CRs, and rate regions for both downlink and uplink scenarios. The main contributions are summarized as follows.
@@
  • For downlink ISAC, we propose three continuous DFSAC beamforming designs: i) a communications-centric (C-C) design that maximizes the CR, ii) a sensing-centric (S-C) design that maximizes the SR, and iii) a Pareto-optimal design that characterizes the Pareto boundary of the rate region. For both C-C and S-C designs, we derive closed-form expressions for SRs and CRs, and examine the asymptotic performance for the infinitely larger CAPAs to gain more physical or geometric insights. For the Pareto-optimal design, we propose a signal subspace-based approach to derive closed-form solutions for the optimal beamformer and the achievable rate region.
@@
  • We present simulation results to demonstrate that: i) the proposed subspace approach for solving the optimal Pareto beamforming offers significantly lower computational complexity compared to the existing Fourier-based method; ii) The achievable SR-CR rate regions of SPDA-based ISAC and CAPA-based FDSAC systems are entirely encompassed within those of the CAPA-based ISAC system in both downlink and uplink scenarios.
@@ II-B1 Communication Model
For LoS scenarios111 While this work primarily focuses on LoS channels to establish an analytically tractable framework for CAPA-based ISAC and to facilitate theoretical investigations into fundamental performance limits, the proposed CAPA-based methodology is general and applicable to arbitrary channel models. In our recent work [40], we have developed a CAPA-based multipath fading channel model, some results of which are directly applicable to ISAC scenarios. These developments provide a solid foundation for future extensions. However, a comprehensive analysis of more complex multipath fading environments is left for future work., the channel response $h_{\rm{c}}(\mathbf{u},{\mathbf{t}})$ is expressed as follows [28]:
@@
where $\sqrt{\frac{\lvert{\mathbf{e}}_{\rm{t}}^{\mathsf{T}}(\mathbf{u}-\mathbf{t})\rvert}{\lVert\mathbf{u}-\mathbf{t}\rVert}}$ accounts for the effective aperture loss caused by misalignment between the signal propagation direction and the array’s normal direction. Moreover, ${\mathbf{e}}_{y}=[0,1,0]^{\mathsf{T}}$ denotes the normal vector to the transmit CAPA, and $g(\cdot,\cdot)$ is the scalar Green’s function that models the effects of free-space EM propagation. By [28], we have
$g(\mathbf{u},\mathbf{t})=-{\rm{j}}\eta k_{0}\frac{{\rm{e}}^{-{\rm{j}}k_{0}\lVert\mathbf{u}-\mathbf{t}\rVert}}{\sqrt{4\pi}\lVert\mathbf{u}-\mathbf{t}\rVert}\left(1+\frac{{\rm{j}}k_{0}^{-1}}{\lVert\mathbf{u}-\mathbf{t}\rVert}-\frac{k_{0}^{-2}}{\lVert\mathbf{u}-\mathbf{t}\rVert^{2}}\right),$ (9)
@@
$g(\mathbf{u},\mathbf{t})=-{\rm{j}}\eta k_{0}\frac{{\rm{e}}^{-{\rm{j}}k_{0}\lVert\mathbf{u}-\mathbf{t}\rVert}}{\sqrt{4\pi}\lVert\mathbf{u}-\mathbf{t}\rVert}.$ (10)
@@ II-B2 Sensing Model
where $h_{\rm{s}}(\mathbf{r},{\mathbf{t}})$ represents the target response, and $n_{{\rm{s}},\ell}(\mathbf{r})$ is zero-mean additive noise satisfying
@@
In this work, assuming that the target’s location is known or has been estimated via a separate process [43, 44, 41, 45], we focus on estimating the RCS $\beta_{\rm{s}}$. The RCS encodes critical information about the target’s physical and EM characteristics, which plays a key role in target recognition, surface material classification, and other high-resolution sensing applications [46, 39]. Under this scenario, the sensing task aims to extract the environmental information contained in $\beta_{\rm{s}}$ from the received echo signals $\{y_{{\rm{s}},\ell}(\mathbf{r})\}_{\ell=1}^{L}$, given knowledge of ${\mathbf{x}}({\mathbf{t}})$. According to the Swerling-I model [47], the RCS $\beta_{\rm{s}}$ is assumed to remain relatively constant across pulses and has an a prior Rayleigh-distributed amplitude. Under this model, $\beta_{\rm{s}}$ follows a complex Gaussian distribution, i.e., $\beta_{\rm{s}}\sim{\mathcal{CN}}(0,\alpha_{\rm{s}})$, where $\alpha_{\rm{s}}>0$ is the arithmetic mean of all RCS values of the reflecting object. This parameter represents the average reflection strength [47].
@@ II-C Uplink ISAC
$x_{{\rm{c}},\ell}({\mathbf{u}})=\sqrt{\frac{P_{\rm{c}}}{\lvert{\mathcal{A}}\rvert}}s_{{\rm{c}},\ell},$ (20)
@@ III-A2 Performance Metric for Sensing
The next task is to recover $\beta_{\rm{s}}$ from $\hat{\mathbf{y}}_{{\rm{s}}}$ given ${\mathbf{s}}$. The information-theoretic limits of this sensing task are measured by the sensing mutual information (MI), which quantifies the MI between $\hat{\mathbf{y}}_{{\rm{s}}}$ and $\beta_{\rm{s}}$, conditioned on the ISAC signal ${\mathbf{s}}$ [39, 45, 48]. To evaluate sensing performance, we use the SR as the performance metric, which is defined as the sensing MI per unit time [39]. Since each DFSAC symbol lasts one unit of time, the SR is expressed as follows:
@@
Notably, the following theorem reveals the connection between the SR and both the mean-squared error (MSE) and the Bayesian Cramér-Rao bound (BCRB) under the considered sensing model.
Theorem 1.
Under the considered sensing model, maximizing the achievable SR is equivalent to minimizing the MSE and BCRB in estimating the RCS.
Proof:
Please refer to Appendix -A for more details. ∎
@@ III-B2 Performance of Communications
Please refer to Appendix -B for more details. ∎
To further elucidate the properties of CAPAs in ISAC and gain more physical or geometric insights, we examine the asymptotic performance by considering the infinitely larger CAPAs, i.e., $L_{x},L_{z}\rightarrow\infty$. The following corollary provides the asymptotic CR under the C-C design.
Corollary 1.
As $L_{x},L_{z}\rightarrow\infty$, the downlink CR achieved by the C-C design satisfies
$\underset{L_{x},L_{z}\rightarrow\infty}{\lim}\mathcal{R}_{\mathrm{d},\mathrm{c}}^{\mathrm{c}}=\log_{2}\left(1+\frac{\overline{\gamma}_{\mathrm{c}}\eta^{2}k_{0}^{2}}{2}\right).$ (41)
Proof:
This can be obtained based on $\underset{L_{x},L_{z}\rightarrow\infty}{\lim}g_{\mathrm{d}}=\frac{\eta^{2}k_{0}^{2}}{4\pi}\times 4\frac{\pi}{2}=\frac{\eta^{2}k_{0}^{2}}{2}$. ∎
Remark 2.
The results of Corollary 1 suggest that the CR will converge to a finite value as the aperture size grows, which is intuitively reasonable from the perspective of energy conservation.
@@ III-B3 Performance of Sensing
$\begin{split}\rho_{\mathrm{d}}=&\frac{\pi^{2}L_{x}L_{z}}{4N^{2}}\sum\nolimits_{n=1}^{N}{\sum\nolimits_{n^{\prime}=1}^{N}{\sqrt{\left(1-\xi_{n}^{2}\right)(1-\xi_{n^{\prime}}^{2})}}}\\ &\times\tilde{h}_{\mathrm{c}}^{*}\Big{(}\frac{\xi_{n}+1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)}\tilde{h}_{\mathrm{s}}\Big{(}\frac{\xi_{n}+1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)},\end{split}$ (45)
@@
$\tilde{h}_{k}(x,z)\triangleq\frac{\mathrm{j}\eta k_{0}\sqrt{\frac{r_{k}\Psi_{k}}{4\pi}}\mathrm{e}^{-\mathrm{j}k_{0}(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{\frac{1}{2}}}}{(x^{2}\!+\!z^{2}\!-\!2r_{k}\!\left(\Phi_{k}x\!+\!\Theta_{k}z\right)\!+\!r_{k}^{2})^{\frac{3}{4}}}$
@@
The term $\int_{\mathcal{A}_{\mathrm{r}}}{\left|a_{\mathrm{r}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}=g_{\mathrm{r}}$ can be computed by following steps similar to those in Appendix -B. Additionally, the term $\rho_{\mathrm{d}}$ can be expressed as follows:
$\rho_{\mathrm{d}}=\int_{-{L_{z}}/{2}}^{{L_{z}}/{2}}{\int_{0}^{L_{x}}{\tilde{h}_{\mathrm{c}}^{*}(x,z)\tilde{h}_{\mathrm{s}}(x,z)\mathrm{d}x\mathrm{d}z}}.$ (46)
@@
We further consider the asymptotic SR for infinitely large arrays. While $\underset{L_{x},L_{z}\rightarrow\infty}{\lim}g_{\mathrm{r}}=\frac{\eta^{2}k_{0}^{2}}{2}$ can be easily obtained, determining $\underset{L_{x},L_{z}\rightarrow\infty}{\lim}\rho_{\mathrm{d}}$ is more challenging. Although one might intuitively expect it to vanish, recent work in [50] employs the stationary phase method to derive an expression for $\underset{L_{x},L_{z}\rightarrow\infty}{\lim}\rho_{\mathrm{d}}$, demonstrating that it converges to a finite value close but not equal to zero. Due to the complexity of this expression, it is omitted here; however, numerical simulations illustrating the asymptotic SR behavior are presented in Section V. These results suggest that as the aperture size increases, the SR achieved by the C-C design asymptotically converges to a finite value larger than zero.